Business Weekly conducted a survey of graduates from 30 top MBA programs. On the basis of the survey, assume the mean annual salary for graduates 10 years after graduation is $160,000 dollars. Assume the standard deviation is $41,000. Suppose you take a simple random sample of 82 graduates.

1) What is the standard deviation of the sampling distribution for this situation? Round to four places. Show work and answer using proper notation.

2) Find the probability that a single randomly selected salary has a value between $143,700 and $164,981. Your write up should include all of the following:

  • Define the variable you are using.
  • Use correct notation to label all quantities.
  • Show the calculations for the relevant z-scores using proper notation.
  • State or show how you found your final probability. You can use tables or StatCrunch, just specify what you did.
  • Write your answer using full and correct probability notation. Round to four places.

3) Find the probability that a random sample of size n=82n=82 has a mean value between $143,700 and $164,981.

  • Define the variable you are using.
  • Use correct notation to label all quantities.
  • Show the calculations for the relevant z-scores using proper notation.
  • State or show how you found your final probability. You can use tables or StatCrunch, just specify what you did.
  • Write your answer using full and correct probability notation. Round to four places.

Write your answers in complete sentence form.
Answer #1

1)  standard deviation of the sampling distribution=σ/√n = 4527.6926

2)

we need to calculate probability for ,                                          
143700   ≤ X ≤    164981                                  
X1 =    143700   ,    X2 =   164981                          
                                          
Z1 =   (X1 - µ )/(σ/√n) = (   143700   -   160000   ) / (   41000   / √   1   ) =   -0.40  
Z2 =   (X2 - µ )/(σ/√n) = (   164981   -   160000   ) / (   41000   / √   1   ) =   0.12  
                                          
P (   143700   < X <    164981   ) =    P (    -0.40   < Z <    0.12   )       
                                          
= P ( Z <    0.12   ) - P ( Z <   -0.40   ) =    0.548   -    0.345   =    0.2029       (answer)
the probability that a single randomly selected salary has a value between $143,700 and $164,981 is 0.2029

3)

we need to calculate probability for ,                                          
143700   ≤ X ≤    164981                                  
X1 =    143700   ,    X2 =   164981                          
                                          
Z1 =   (X1 - µ )/(σ/√n) = (   143700   -   160000   ) / (   41000   / √   82   ) =   -3.60  
Z2 =   (X2 - µ )/(σ/√n) = (   164981   -   160000   ) / (   41000   / √   82   ) =   1.10  
                                          
P (   143700   < X <    164981   ) =    P (    -3.60   < Z <    1.10   )       
                                          
= P ( Z <    1.10   ) - P ( Z <   -3.60   ) =    0.864   -    0.000   =    0.8642       (answer)
the probability that a random sample of size n=82 has a mean value between $143,700 and $164,981 is 0.8642